OpenStudy (anonymous):

For two functions, a(x) and b(x), a statement is made that a(x) = b(x) at x = 2. What is definitely true about x = 2? Both a(x) and b(x) have a maximum or minimum value at x = 2. Both a(x) and b(x) have the same output value at x = 2. Both a(x) and b(x) cross the x-axis at 2. Both a(x) and b(x) cross the y-axis at 2.

OpenStudy (anonymous):

Well, do they tell you what a(x) or b(x) equal besides each other?

OpenStudy (anonymous):

For instance a(x) = x*1 would mean that Both a(x) and b(x) cross the y- axis at 2 when x = 2

OpenStudy (anonymous):

so is it c?

OpenStudy (anonymous):

i mean D

OpenStudy (anonymous):

It is hard to say without knowing what a(x) or b(x) equals other than each other. We know a(x) = b(x) but do you have anything else one of them might equal?

OpenStudy (anonymous):

If a(x) or b(x) is a quadratic then it is a different story. It seem like you are missing something in your problem. Did you type it in right?

OpenStudy (anonymous):

Never mind. I think they just want to know what is true when x = 2 and that would be C

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

They're asking which one is DEFINITELY true. All of them can be true under the right circumstances. a(x) = b(x) at 2 means a(2) = b(2), meaning they have the same y (or output) at x = 2

OpenStudy (anonymous):

That is what I was getting to at first peachpi but then it stats What is definitely true about x = 2 and what is definitely true is when x = 2 it will be a vertical line at 2 so I figure they want you to chose C

OpenStudy (anonymous):

C was wrong

OpenStudy (anonymous):

we don't know that it's a vertical line. We don't know anything except that two function have the same y at x = 2. The only thing consistent is that they have the same y